OpenStudy (anonymous):

A speeder passes a parked police car at a constant speed of 28.4 m/s. At that instant, the police car starts from rest with a uniform acceleration of 2.93 m/s2. How much time passes before the speeder is overtaken by the police car? Answer in units of s

OpenStudy (jamesj):

Let's set x = 0 where the police car is. What's the equation of motion for the two cars? \[ x_{speeder}(t) = ... what? \] \[ x_{police}(t) = ... what? \]

OpenStudy (jamesj):

the speeder's location xs(t) is xs(t) = 28.4t Now what's the police car's position, xp(t) ?

OpenStudy (anonymous):

(1/2)(2.93)t^2

OpenStudy (jamesj):

Right, exactly. Now when are the two equal?

OpenStudy (anonymous):

set them = to each other?

OpenStudy (anonymous):

when the police car catches the speeder

OpenStudy (jamesj):

Set them equal and find the value of t for which that is the case.

OpenStudy (anonymous):

48.46 and 0

OpenStudy (anonymous):

obviously its not 0

OpenStudy (jamesj):

Right. t = 0 is when the driver first goes past the police car. You want the other answer.

OpenStudy (anonymous):

and it is not correct. I've already tried it previously.

OpenStudy (jamesj):

t = 48 sec isn't right. Check your calculation

OpenStudy (jamesj):

\[ 28.4 t = \frac{2.93}{2} t^2 \]

OpenStudy (jamesj):

\[ t = 0 \] or \[ t = \frac{2 \times 28.4}{2.93} \]

OpenStudy (anonymous):

I made an error somewhere, thank-you.

OpenStudy (anonymous):

2.93? wait

OpenStudy (anonymous):

2 * a a=.586?

OpenStudy (anonymous):

.586^2 ?

OpenStudy (anonymous):

noooo

OpenStudy (anonymous):

1/2 * 2.93?

OpenStudy (jamesj):

The position of the speeder is\[ x_s(t) = 28.4 t \]The position of the police car is \[ x_p(t) = \frac{1}{2} 2.93t^2 \] Hence the positions are equal when \[ 28.4 t = \frac{1}{2}2.93t^2 \] Now solve for t as I've written above. You can factor out a t from both sides if t is not zero, which gives you \[ \frac{1}{2}2.93t = 28.4 \] hence \[ t = \frac{2 \times 28.4}{2.93} \] Now simplify that expression.

OpenStudy (anonymous):

19.39

OpenStudy (jamesj):

Yes, t = 19.39 seconds

OpenStudy (anonymous):

I was making this more difficult by using quadratic equation. Thank you for the help and clarification.